# Direct probing the quantum geometric tensor for bosonic collective excitations

### v1

@@ Abstract
I Introduction
Quantum geometry tensor (QGT) characterizes the intrinsic Hilbert-space structure of quantum states; its real and imaginary parts define the quantum metric tensor (QMT) and Berry curvature, respectively [5, 4, 64, 44, 9]. Direct observation of the QGT and its components is essential for elucidating its impact on condensed matter phenomena. For example, the Berry curvature governs topological phases and transport responses[46, 61, 8, 52], while the quantum metric in electronic systems has emerged as a key ingredient in phenomena ranging from nonlinear Hall effects [10, 60, 45, 11, 14, 32] to flat-band superconductivity [62, 58, 31, 65, 42, 21, 15, 16, 41, 55, 25], anomalous Landau quantization [47, 19, 22, 22], and excitonic shifts [53, 38, 63, 59], among others [34, 43, 12, 1, 26, 37, 6, 30, 35, 39, 56, 68, 7, 40, 29]. While the QMT has been experimentally accessed in artificial platforms [57, 66, 13, 49] and solids via techniques such as angle-resolved photoemission spectroscopy (ARPES), these measurements have so far been confined to electronic bands [23, 24]. Consequently, direct experimental access to the QMT, and thus the full QGT, for bosonic collective excitations remains elusive. Bridging this gap is a crucial step toward understanding the role of quantum geometry in bosonic systems.
To overcome this challenge, we develop a general framework that enables direct access to the full QGT, including both the Berry curvature and quantum metric, of collective excitations via the dynamical structure factor (DSF) [54]. By formulating the DSF and QGT within a unified pseudospin framework, we establish a direct correspondence between measurable scattering intensities and the complete QGT. Although illustrated using two-band (subspace) Hamiltonians, the approach applies generally to multiband systems. As a universal probe of density-fluctuation-driven quasiparticles, the DSF enables a practical protocol, implementable with inelastic x-ray or neutron scattering, to resolve all components of the bosonic QGT across the Brillouin zone. We demonstrate the approach with the twofold quadrupole Weyl phonon in BaPtGe [67, 28] and the nodal-line magnon in Gd [50], revealing clear signatures of quantum geometry.
II Bosonic QGT
@@
The QGT is defined in a $\mathrm{U(1)}$-gauge-invariant form as
@@
where $\pm$” donates the upper/lower band of the two-band system (Details are in the supplementary). Consequently, direct access to the pseudospin vector $\mathbf{d}$ would immediately determine both $\Omega_{n}^{ab}$ and $g^{ab}_{n}$. In contrast to electronic systems, where circularly polarized ARPES can resolve the pseudospin [23, 24], the polarization of the incident probe usually does not directly imprint on the measured response of bosonic excitations. This essential difference rules out an ARPES-like strategy and calls for a fundamentally different approach.
@@
III Phonon DSF and QGT
The DSF characterizes a system’s response to momentum- and energy-resolved probes, such as neutrons or x-rays [27, 2, 3, 18]. As illustrated in Figs. 1(a)-(b), it encodes the spectrum of density fluctuations. The total DSF reads
$S(\mathbf{Q},\omega)_{\text{1p}}=$ $N_{\text{cell}}\sum_{\mathbf{k},\sigma}\frac{1}{\omega_{\mathbf{k},\sigma}}\Big|\sum_{d}\frac{f_{d}(\mathbf{Q})}{\sqrt{m_{d}}}e^{-W_{d}}\{\mathbf{Q}\cdot\xi_{d}(\mathbf{k})\}e^{i\mathbf{Q\cdot r}_{j}}\Big|^{2}\times$ (3)
$\Big\{(\langle n_{\mathbf{k},\sigma}\rangle+1)\delta(\omega-\omega_{\sigma}(\mathbf{k}))\delta(\mathbf{Q-G-k})+$
$\langle n_{\mathbf{k},\sigma}\rangle\delta(\omega+\omega_{\sigma}(\mathbf{k}))\delta(\mathbf{Q-G+k})\Big\}.$
@@ III Phonon DSF and QGT
$S_{\text{coh}}^{(s)}(\mathbf{Q},\omega)\propto|F_{\text{coh}}^{(s)}(\mathbf{Q})|^{2}=\left|\sum_{d}\frac{f_{d}(\mathbf{Q})}{\sqrt{m_{d}}}e^{-W_{d}}\mathbf{Q}\cdot\xi_{d}(\mathbf{k})e^{i\mathbf{Q}\cdot\mathbf{r_{d}}}\right|^{2}.$ (4)
$\mathbf{Q=G+k}$ denotes the total momentum transfer, $f_{d}(\mathbf{Q}),m_{d},W_{d},\mathbf{r_{d}}$ and $\xi_{d}(\mathbf{k})$ represent the atomic form factor, mass of atom, Debye-Waller factor, equilibrium position, and $s^{\text{th}}$ polarization vector of the $d^{\text{th}}$ atom in the unit cell, respectively. For inelastic x-ray scattering, $f_{d}(\mathbf{Q})$ depends on momentum transfer $\mathbf{Q}$, whereas for neutron scattering it is a constant (the scattering length). Equation (4) reveals that the DSF is primarily governed by the factor $\mathbf{Q}\cdot{\xi}_{d}(\mathbf{k})$, which directly links the measured scattering spectra to the underlying microscopic dynamics, including the full phonon polarization structure.
@@
$=\frac{1}{2}\mathbf{|V|}\Big\{1+\cos\theta_{\mathbf{k}}\cos\theta_{\mathbf{G}}+\sin\theta_{\mathbf{k}}\sin\theta_{\mathbf{G}}\cos(\phi_{\mathbf{k}}-\phi_{\mathbf{G}})\Big\},$
where $W^{\pm}(\mathbf{G})=~\sum_{d}\frac{f_{d}(\mathbf{Q})}{\sqrt{m_{d}}}e^{-W_{d}}\mathbf{G}\cdot\xi_{d}^{\pm}e^{i\mathbf{G}\cdot\mathbf{r_{d}}}$ are two components of $\langle W|$. The vectors
@@
IV Phonon QGT in BaPtGe
As an example, we next consider the twofold quadrupole Weyl (TQW) phonon at $\Gamma$ in chiral cubic crystal BaPtGe [67, 28, 20], as shown in Figs. 2(a)-(c). We begin with a simplified $k\cdot p$ model and subsequently present the corresponding first-principles results. The effective model for the TQW is described by the Hamiltonian:
@@ IV Phonon QGT in BaPtGe
$W^{\pm}(\mathbf{Q}\approx\mathbf{G})=\frac{2\pi}{a}\frac{b_{\text{coh}}\cdot e^{-W}}{\sqrt{12m}}\cdot\{g_{x}A_{1}+g_{y}\omega^{\pm 1}A_{2}+g_{z}\omega^{\mp 1}A_{3}\}.$ (11)
@@
Figure 2(d) shows the heat-map distribution of $\cos\langle\mathbf{d},\mathbf{V}\rangle=\sin\theta_{\mathbf{k}}\sin\phi_{\mathbf{k}}$ from the $k\cdot p$ model for $\mathbf{G}^{V_{y}}=(3,1,0)$, plotted on the $\alpha$ plane shown in Fig. 2(b). Using these distributions at different $\mathbf{Q}$, we extract the pseudospin components $d_{x,y,z}$, as shown in Fig. 2(e) (Details are in the Supplementary Material). Once $\mathbf{d}(\mathbf{k})$ is determined, the full quantum geometric tensor can be directly evaluated, such as the $g^{xx}$ component shown in Fig. 2(f).
In this two-band model, the quantum metric $g^{ab}$ encodes the same geometric volume as the Berry curvature $\Omega^{ab}$, so that the Chern number and the determinant of the quantum metric coincide with the Berry curvature, i.e., $\Omega^{ab}=2\sqrt{\text{det}(g^{ab})}$, as shown in Fig. 2(g) in polar coordinates [36, 42, 33, 48].
V Symmetry analysis and DFT calculations on BaPtGe
The QGT for TQW in BaPtGe is constrained by the little co-group at the $\Gamma$ point, which includes the twofold rotations $C_{2x,2y,2z}$, the threefold rotation $C_{3}^{[111]}$, and time-reversal symmetry $\mathcal{T}$. The corresponding transformation rules are summarized in the Supplementary Material. On the high-symmetry planes, these symmetries impose the following constraints:
@@ VI Magnon DSF in Gadolinium
Elemental gadolinium (Gd) has been reported to host both node-line magnons (along HKH-K) and node-plane magnons (on the $q_{z}=\pi$ plane) [50, 51], as shown in Figs. 4 (a)-(b). Near the magnon Dirac cone at the $K$ point, inelastic neutron scattering (INS) exhibits an anisotropic intensity pattern, which is well captured by linear spin-wave theory (LSWT) and the two-band $k\cdot p$ model below:
@@
where $\phi_{\mathbf{q}}=\arctan d_{y}(\mathbf{q})/d_{x}(\mathbf{q})$ and $\mathbf{d(q)}=\begin{pmatrix}\cos\phi_{\mathbf{q}}&\sin\phi_{\mathbf{q}}&0\end{pmatrix}$.
@@
Thus, the DSF of the node-line magnon, shown in Fig. 4(c), exhibits a $\cos\phi_{\mathbf{q}}$-like dependence on the constant-energy surface in momentum space, while the factor $(-1)^{g_{z}}$ leads to alternating intensity between the upper and lower bands as the momentum crosses the Brillouin zone boundary along $g_{z}$ direction, matching with the observation in experiments [50].
@@
Figures 4(e)–(f) display $g^{yy}$ on a constant-energy surface in momentum space and along $\Gamma-K$, respectively. Along $\Gamma-K$, $g^{yy}_{\pm}=\frac{q^{2}_{x}}{4|\mathbf{q}|^{4}}\propto\frac{1}{q_{x}^{2}}$ leading to a pronounced enhancement near the $K$ point and a vanishing value along the $q_{z}$ axis.
@@
VII Conclusion
In summary, we establish the dynamical structure factor as a direct experimental probe of both the Berry curvature and quantum metric, enabling complete access to the quantum geometric tensor across the Brillouin zone. Validated in twofold quadruple Weyl phonons and node-line magnons, our approach unifies experimentally measurable scattering intensities with quantum geometry through a pseudospin framework and is readily extendable to multi-band systems with well-defined pseudospin degrees of freedom. By directly linking the DSF to quantum geometry, our work enables experimental access not only to spectral properties but also to geometric and topological responses, including sensitivity to perturbations, topology-induced mode dynamics, enhanced transport, and Chern numbers, thereby establishing a general paradigm for exploring quantum-geometric phenomena in bosonic quasiparticles and paving the way for understanding its role in condensed matter systems.
VIII Note added
We note a related work on the resolution of topology and geometry from momentum-resolved spectroscopies [17].
IX Acknowledgements
We acknowledge the helpful discussion with H. Miao. T. Zhang and C. Wu acknowledge the support from National Key R&D Project (Grant Nos. 2023YFA1407400 and 2024YFA1409200), and the National Natural Science Foundation of China (Grant Nos. 12374165 and 12447101). T.O. acknowledge support from JSPS KAKENHI (No. JP23H04865, No. 23K22487), MEXT, Japan. S.M. is supported by JSPS KAKENHI Grant Nos. JP22H00108 and JP24H02231, and also by MEXT Initiative to Establish Next-generation Novel Integrated Circuits Centers (X-NICS) Grant No. JPJ011438.

### v2

@@ Abstract
Introduction.–Quantum geometry tensor (QGT) characterizes the intrinsic Hilbert-space structure of quantum states; its real and imaginary parts define the quantum metric tensor (QMT) and Berry curvature, respectively Berry (1984, 1989); Yu et al. (2025); Provost and Vallee (1980); Cheng (2013). Direct observation of the QGT and its components is essential for elucidating its impact on condensed matter phenomena. For example, the Berry curvature governs topological phases and transport responsesThouless et al. (1982); Xiao et al. (2010); Chang and Niu (2008); Sodemann and Fu (2015), while the quantum metric in electronic systems has emerged as a key ingredient in phenomena ranging from nonlinear Hall effects Gao et al. (2023); Wang et al. (2023); Qiang et al. (2025); Gao et al. (2014); Han et al. (2024); Liu et al. (2021) to flat-band superconductivity F. Xie, Z. Song, B. Lian, and B. A. Bernevig (2020); H. Tian, X. Gao, Y. Zhang, S. Che, T. Xu, P. Cheung, K. Watanabe, T. Taniguchi, M. Randeria, F. Zhang, et al. (2023); L. Liang, T. I. Vanhala, S. Peotta, T. Siro, A. Harju, and P. Törmä (2017); J. Yu, C. J. Ciccarino, R. Bianco, I. Errea, P. Narang, and B. A. Bernevig (2024); S. Peotta and P. Törmä (2015); A. Julku, S. Peotta, T. I. Vanhala, D. Kim, and P. Törmä (2016); J. Herzog-Arbeitman, V. Peri, F. Schindler, S. D. Huber, and B. A. Bernevig (2022); J. S. Hofmann, E. Berg, and D. Chowdhury (2023); 45; Z. Sun, R. Yu, S. A. Chen, J. Hu, and K. T. Law (2025); T. Kitamura, A. Daido, and Y. Yanase (2025), anomalous Landau quantization Rhim et al. (2020); Hwang et al. (2021); Jung et al. (2024, 2024), and excitonic shifts Srivastava and Imamoğlu (2015); Paiva et al. (2024); Ying and Law (2024); Verma et al. (2024), among others Neupert et al. (2013); Piéchon et al. (2016); Gao et al. (2015); Albert et al. (2016); Kolodrubetz et al. (2017); Ozawa (2018); Bleu et al. (2018); Li et al. (2025); Ouyang et al. (2025); Paul (2024); Tam et al. (2024); Zhou (2024); Bzdušek (2025); Pellitteri et al. (2025); Li and Liu (2025). While the QMT has been experimentally accessed in artificial platforms Tan et al. (2019); Yu et al. (2020); Gianfrate et al. (2020); Sala et al. (2025) and solids via techniques such as angle-resolved photoemission spectroscopy (ARPES), these measurements have so far been confined to electronic bands Kang et al. (2025); Kim et al. (2025). Consequently, direct experimental access to the QMT, and thus the full QGT, for bosonic collective excitations remains elusive. Bridging this gap is a crucial step toward understanding the role of quantum geometry in bosonic systems.
To overcome this challenge, we develop a general framework that enables direct access to the full QGT, including both the Berry curvature and quantum metric, of collective excitations via the dynamical structure factor (DSF) Sturm (1993). By formulating the DSF and QGT within a unified pseudospin framework, we establish a direct correspondence between measurable scattering intensities and the complete QGT. Although illustrated using two-band (subspace) Hamiltonians, the approach applies generally to multiband systems. As a universal probe of density-fluctuation-driven quasiparticles, the DSF enables a practical protocol, implementable with inelastic x-ray or neutron scattering, to resolve all components of the bosonic QGT across the Brillouin zone. We demonstrate the approach with the twofold quadrupole Weyl phonon in BaPtGe Zhang et al. (2020); Li et al. (2021) and the nodal-line magnon in Gd Scheie et al. (2022a), revealing clear signatures of quantum geometry.
@@
Bosonic QGT.–The QGT is defined in a $\mathrm{U(1)}$-gauge-invariant form as
@@
where $\pm$” donates the upper/lower band of the two-band system (Details are in the supplementary). Consequently, direct access to the pseudospin vector $\mathbf{d}$ would immediately determine both $\Omega_{n}^{ab}$ and $g^{ab}_{n}$. In contrast to electronic systems, where circularly polarized ARPES can resolve the pseudospin Kang et al. (2025); Kim et al. (2025), the polarization of the incident probe usually does not directly imprint on the measured response of bosonic excitations. This essential difference rules out an ARPES-like strategy and calls for a fundamentally different approach.
@@
Phonon DSF and QGT.–The DSF characterizes a system’s response to momentum- and energy-resolved probes, such as neutrons or x-rays Kotani and Shin (2001); Ament et al. (2011); Benedek and Toennies (1994); Hudson (2006). As illustrated in Figs. 1(a)-(b), it encodes the spectrum of density fluctuations. The total DSF reads
$S(\mathbf{Q},\omega)_{\text{1p}}=$ $\sum_{\mathbf{k},\sigma}\frac{N_{\text{cell}}}{\omega_{\mathbf{k},\sigma}}\Big|\sum_{d}\frac{f_{d}(\mathbf{Q})}{\sqrt{m_{d}}}e^{-W_{d}}\{\mathbf{Q}\cdot\xi_{d}(\mathbf{k})\}e^{i\mathbf{Q\cdot r}_{j}}\Big|^{2}$ (3)
$\times\Big\{(\langle n_{\mathbf{k},\sigma}\rangle+1)\delta(\omega-\omega_{\sigma}(\mathbf{k}))\delta(\mathbf{Q-G-k})$
$+\langle n_{\mathbf{k},\sigma}\rangle\delta(\omega+\omega_{\sigma}(\mathbf{k}))\delta(\mathbf{Q-G+k})\Big\}.$
@@
$S_{\text{coh}}^{(s)}(\mathbf{Q},\omega)$ $\propto|F_{\text{coh}}^{(s)}(\mathbf{Q})|^{2}$ (4)
$=\left|\sum_{d}\frac{f_{d}(\mathbf{Q})}{\sqrt{m_{d}}}e^{-W_{d}}\mathbf{Q}\cdot\xi_{d}(\mathbf{k})e^{i\mathbf{Q}\cdot\mathbf{r_{d}}}\right|^{2}.$
$\mathbf{Q=G+k}$ denotes the total momentum transfer, $f_{d}(\mathbf{Q}),m_{d},W_{d},\mathbf{r_{d}}$ and $\xi_{d}(\mathbf{k})$ represent the atomic form factor, mass of atom, Debye-Waller factor, equilibrium position, and $s^{\text{th}}$ polarization vector of the $d^{\text{th}}$ atom in the unit cell, respectively. For inelastic x-ray scattering, $f_{d}(\mathbf{Q})$ depends on momentum transfer $\mathbf{Q}$, whereas for neutron scattering it is a constant (the scattering length). Eq. (4) reveals that the DSF is primarily governed by the factor $\mathbf{Q}\cdot{\xi}_{d}(\mathbf{k})$, which directly links the measured scattering spectra to the underlying microscopic dynamics, including the full phonon polarization structure.
@@
$=\frac{1}{2}\mathbf{|V|}\Big\{1+\cos\theta_{\mathbf{k}}\cos\theta_{\mathbf{G}}$
$+\sin\theta_{\mathbf{k}}\sin\theta_{\mathbf{G}}\cos(\phi_{\mathbf{k}}-\phi_{\mathbf{G}})\Big\},$
where $W^{\pm}(\mathbf{G})=\penalty 10000\ \sum_{d}\frac{f_{d}(\mathbf{Q})}{\sqrt{m_{d}}}e^{-W_{d}}\mathbf{G}\cdot\xi_{d}^{\pm}e^{i\mathbf{G}\cdot\mathbf{r_{d}}}$ are two components of $\langle W|$. The vectors
@@
Phonon QGT in BaPtGe.–As an example, we next consider the twofold quadrupole Weyl (TQW) phonon at $\Gamma$ in chiral cubic crystal BaPtGe Zhang et al. (2020); Li et al. (2021); Jin et al. (2022), as shown in Figs. 2(a)-(c). We begin with a simplified $k\cdot p$ model and subsequently present the corresponding first-principles results. The effective model for the TQW is described by the Hamiltonian:
@@
$W^{\pm}(\mathbf{Q}\approx\mathbf{G})$ $=\frac{2\pi}{a}\frac{b_{\text{coh}}\cdot e^{-W}}{\sqrt{12m}}\cdot\{g_{x}A_{1}$ (11)
$+g_{y}\omega^{\pm 1}A_{2}+g_{z}\omega^{\mp 1}A_{3}\}.$
@@
Fig. 2(d) illustrates the heat-map distribution of $\cos\langle\mathbf{d},\mathbf{V}\rangle=\sin\theta_{\mathbf{k}}\sin\phi_{\mathbf{k}}$ from the $k\cdot p$ model for $\mathbf{G}^{V_{y}}=(3,1,0)$, plotted on the $\alpha$ plane shown in Fig. 2(b). Using these distributions at different $\mathbf{Q}$, we extract the pseudospin components $d_{x,y,z}$, as shown in Fig. 2(e) (Details are in the Supplementary Material). Once $\mathbf{d}(\mathbf{k})$ is determined, the full quantum geometric tensor can be directly evaluated, such as the $g^{xx}$ component shown in Fig. 2(f).
In this two-band model, the quantum metric $g^{ab}$ encodes the same geometric volume as the Berry curvature $\Omega^{ab}$, so that the Chern number and the determinant of the quantum metric coincide with the Berry curvature, i.e., $\Omega^{ab}=2\sqrt{\text{det}(g^{ab})}$, as shown in Fig. 2(g) in polar coordinates Ozawa and Mera (2021); Peotta and Törmä (2015); Liu et al. (2025); Roy (2014).
Symmetry analysis and DFT calculations on BaPtGe. The QGT for TQW in BaPtGe is constrained by the little co-group at the $\Gamma$ point, which includes the twofold rotations $C_{2x,2y,2z}$, the threefold rotation $C_{3}^{[111]}$, and time-reversal symmetry $\mathcal{T}$. The corresponding transformation rules are summarized in the Supplementary Material. On the high-symmetry planes, these symmetries impose the following constraints:
@@
Magnon DSF in Gadolinium.–Elemental gadolinium (Gd) has been reported to host both node-line magnons (along HKH-K) and node-plane magnons (on the $q_{z}=\pi$ plane) Scheie et al. (2022a, b), as shown in Figs. 4 (a)-(b). Near the magnon Dirac cone at the $K$ point, inelastic neutron scattering (INS) exhibits an anisotropic intensity pattern, which is well captured by linear spin-wave theory (LSWT) and the two-band $k\cdot p$ model below:
@@
where $\phi_{\mathbf{q}}=\arctan\frac{d_{y}(\mathbf{q})}{d_{x}(\mathbf{q})}$ and $\mathbf{d(q)}=\begin{pmatrix}\cos\phi_{\mathbf{q}}&\sin\phi_{\mathbf{q}}&0\end{pmatrix}$.
@@
Thus, the DSF of the node-line magnon, shown in Fig. 4(c), exhibits a $\cos\phi_{\mathbf{q}}$-like dependence on the constant-energy surface in momentum space, while the factor $(-1)^{g_{z}}$ leads to alternating intensity between the upper and lower bands as the momentum crosses the Brillouin zone boundary along $g_{z}$ direction, matching with the observation in experiments Scheie et al. (2022a).
@@
Figs. 4(e)–(f) display $g^{yy}$ on a constant-energy surface in momentum space and along $\Gamma-K$, respectively. Along $\Gamma-K$, $g^{yy}_{\pm}=\frac{q^{2}_{x}}{4|\mathbf{q}|^{4}}\propto\frac{1}{q_{x}^{2}}$ leading to a pronounced enhancement near the $K$ point and a vanishing value along the $q_{z}$ axis.
@@
Conclusions.–In summary, we establish the dynamical structure factor as a direct experimental probe of both the Berry curvature and quantum metric, enabling complete access to the quantum geometric tensor across the Brillouin zone. Validated in twofold quadruple Weyl phonons and node-line magnons, our approach unifies experimentally measurable scattering intensities with quantum geometry through a pseudospin framework and is readily extendable to multi-band systems with well-defined pseudospin degrees of freedom. By directly linking the DSF to quantum geometry, our work enables experimental access not only to spectral properties but also to geometric and topological responses, including sensitivity to perturbations, topology-induced mode dynamics, enhanced transport, and Chern numbers, thereby establishing a general paradigm for exploring quantum-geometric phenomena in bosonic quasiparticles and paving the way for understanding its role in condensed matter systems.
Note added.–We note a related work on the resolution of topology and geometry from momentum-resolved spectroscopies Huang and Fang (2026).
Acknowledgements.–We acknowledge the helpful discussion with H. Miao. T. Zhang and C. Wu acknowledge the support from National Key R&D Project (Grant Nos. 2023YFA1407400 and 2024YFA1409200), and the National Natural Science Foundation of China (Grant Nos. 12374165 and 12447101). T.O. acknowledge support from JSPS KAKENHI (No. JP23H04865, No. 23K22487), MEXT, Japan. S.M. is supported by JSPS KAKENHI Grant Nos. JP22H00108 and JP24H02231, and also by MEXT Initiative to Establish Next-generation Novel Integrated Circuits Centers (X-NICS) Grant No. JPJ011438.